# A Simple Framework to Accelerate Multilingual Language Model for Monolingual Text Generation

### v1

@@ Abstract
Recent advancements in large language models have facilitated the execution of complex language tasks, not only in English but also in non-English languages. However, the tokenizers of most language models, such as Llama, trained on English-centric corpora, tend to excessively fragment tokens in non-English languages. This issue is especially pronounced in non-roman alphabetic languages, which are often divided at a character or even Unicode level, leading to slower text generation. To address this, our study introduces a novel framework designed to expedite text generation in these languages. This framework predicts larger linguistic units than those of conventional multilingual tokenizers and is specifically tailored to the target language, thereby reducing the number of decoding steps required. Our empirical results demonstrate that the proposed framework increases the generation speed by a factor of 1.9 compared to standard decoding while maintaining the performance of a pre-trained multilingual model on monolingual tasks.
@@ 1 Introduction
The advent of large language models (LLMs) (OpenAI, 2023; Touvron et al., 2023; Antropic, 2023) has revolutionized the field of generative artificial intelligence, demonstrating remarkable capabilities to respond across diverse tasks and languages (Eloundou et al., 2023; Solaiman et al., 2023). These models, despite being predominantly trained on English data, exhibit significant multilingual proficiency (Bandarkar et al., 2023).
@@
To address these challenges, previous studies Wang et al. (2019b); Rust et al. (2021); Cui et al. (2023) have proposed replacing or augmenting the existing vocabulary of pre-trained multilingual models with language-specific vocabularies to more effectively encode monolingual text corpora. Specifically, Rust et al. (2021) improved mBERT Devlin et al. (2019) by replacing its tokenizer with a monolingual one and incorporating an additional 100,000 pre-training steps. On the other hand, Cui et al. (2023) enhanced Llama (Touvron et al., 2023) by expanding the Chinese vocabulary and further pre-training it on a 120GB text corpus that includes Chinese texts. However, this approach requires an extensive pre-training phase with a substantial amount of data.
@@
In response to these challenges, our research introduces MuMo, a novel framework designed to accelerate Multilingual language models for targeted Monolingual text generation, particularly in non-alphabetic languages. MuMo incorporates a new vocabulary of a target language into the output layer, also known as the Language Model (LM) head, and predicts the next token from this expanded vocabulary. This approach requires training only the extended portion of the output layer and specific layers of the feed-forward network. Importantly, MuMo eliminates the need for extensive text corpora or a draft model, requiring only a modest corpus of the target language, approximately 365M tokens in size.
Empirical results across summarization, and translation tasks in Korean and Japanese demonstrate that MuMo significantly accelerates text generation, achieving over a 1.9-fold increase in speed without compromising output quality.
The contributions of our research are as follows:
  • We propose MuMo, a framework that accelerates the decoding speed of non-Roman alphabetic languages without compromising the performance of the language model.
  • Our study shows a novel approach to address the issue of slow inference speed of non-English languages, a problem that arises due to the excessive fragmentation in pre-trained models that are primarily English-centric.
  • It can be effectively employed even in the absence of extensive computing resources or large monolingual text corpora.
2 Background
Language Modeling
The goal of auto-regressive language modeling is to create a probability distribution, denoted as $p_{\theta}(\mathbf{x})$, where $\theta$ represents the model parameter. This distribution is established over a sequence of tokens, represented as $\mathbf{x}=[x_{1},\ldots,x_{T}]$. These tokens are derived from a pre-established vocabulary, denoted as $\mathcal{V}$, where each token $x_{t}$ is a member of $\mathcal{V}$. A conventional approach involves maximizing the log-likelihood of a sequence:
$\max_{\theta}\mathcal{L}_{\text{MLE}}(p_{\theta},\mathbf{x})=\sum_{t=1}^{T}\log p_{\theta}(x_{t}|\mathbf{x}_{<t}).$ (1)
Tokenization in LM
Tokenization is the process of segmenting text into subword chunks (Kudo, 2018; Sennrich et al., 2016). The process can be started with a base vocabulary composed solely of characters, then, new vocabulary items are subsequently added by recursively merging existing ones based on the frequency statistics within a training corpus. However, in a multilingual setup, even a base vocabulary composed solely of characters can be extensive, with over 130K items (Wang et al., 2019a).
Byte-level Subword Tokenization
Radford et al. (2019) proposed using bytes as the basic units, with only 256 tokens required to form subwords. This Byte-level Byte Pair Encoding (BBPE) does not suffer from out-of-vocabulary problem and has become the standard in most pre-trained models (Brown et al., 2020; Muennighoff et al., 2022; Zhang et al., 2022b; Scao et al., 2022). However, BBPE often produces longer sequence length when handling multilingual scripts compared to character-based BPE (Wang et al., 2019a; Petrov et al., 2023).
Lang Word Multilingual Tokens Ko 서울 (“서”, “\0xec”, “\0xb8”, “\0x9a”) Ja 発売 (“発”, “\0xe5”, “\0xa3”, “\0xb2
@@ 2 Background
Table 1: Examples of the tokenization results. These examples are tokenized by the Llama tokenizer. The target monolingual word are excessively segmented into byte units, when a suitable match is not found in the multilingual vocabulary.
Exccessive Fragmentation
We explore the issue of excessive fragmentation (Ahia et al., 2023; Petrov et al., 2023) in multilingual setups, a common challenge in multilingual tokenizers (Rust et al., 2021; Zhang et al., 2022a; Ahia et al., 2023; Petrov et al., 2023). The problem occurs when a tokenizer produces too many tokens due to significant differences between input texts and the pre-trained corpus (Rust et al., 2021; Hong et al., 2021).
The difficulty of integrating numerous languages into one model while balancing vocabulary composition contributes to this issue (Wang et al., 2019b). As the number of vocabulary items increases, model parameters expand exponentially, impacting training efficiency (Liang et al., 2023).
To investigate this, we compared the tokenization outputs of several multilingual model tokenizers: Bloomz (Scao et al., 2022), Llama-2 (Touvron et al., 2023), and GPT-3.5/4 (OpenAI, 2023). We used the FLoRes-200 (Goyal et al., 2022) dataset, which includes Wikipedia articles in 200 languages. Table 2 indicates that multilingual tokenizers generate more tokens for Japanese and Korean languages than their monolingual counterparts.
@@ 3 Proposed Framework
Figure 2: Illustration of a single step prediction with MuMo. Initially, the MuMo LM Head $f_{\text{mumo}}$ selects the top 6 candidates. Then, the pre-trained multilingual model verifies the feasibility of the candidates. Among the modules in MuMo, the Target Monolingual LM head (the Korean LM Head in the figure) is only trained.
We propose a framework named MuMo to accelerate the inference speed of a pre-trained multilingual LM for a non-alphabetic monolingual language via a given small monolingual dataset. In the section, we introduce 1) the model architecture, 2) the fine-tuning process on a small monolingual dataset, and 3) the inference process of the proposed framework.
Model # of Vocabulary en ja ko Bloomz 250K 26.0 46.5 72.4 Llama-2 32K 29.5 66.3 94.6 GPT-3.5 / 4 100K 25.9 58.6 61.3 Mono-Ja 30K - 28.1 - Mono-Ko 32K - - 32.3
Table 2: Average number of tokens after tokenization. We evaluate tokenization efficiency on the FLoRes-200 dataset for English, Japanese, and Korean Languages. Mono-Ja, and Mono-Ko indicate the tokenizer of pre-trained monolingual model for Japanese and Korean respectively.
@@ 3.1 Model Architecture
We consider a setting in which a pre-trained multilingual model $f_{\text{multi}}$ is given. The model consists of 1) Transformer layers that consist of attention and feed-forward network, and 2) an output embedding layer called language model (LM) head. We denote $\mathcal{V}_{\text{multi}}$ as the multilingual vocabulary set of the model. objective, as $\mathcal{L}_{\text{MLE}}(p_{\text{multi}},\mathbf{x})=\sum_{t=1}^{T}\log p_{\text{multi}}(x_{t}|\mathbf{x}_{<t})$
@@
The primary concept involves modifying pre-trained representations to predict a singular token unit within a target monolingual vocabulary $\mathcal{V}_{\text{mono}}$. We refer to the Target Monolingual LM head as $f_{\text{mono}}$, which is composed of two main components: a feed-forward network (FFN) and an output linear layer, represented as $g_{\text{mono}}:\mathbb{R}^{d_{{}_{\text{mono}}}}\rightarrow\mathbb{R}^{|\mathcal{V}_{\text{mono}}|}$:
$\text{FFN}(h)=q(W_{1}^{\top}h)W_{2}\in\mathbb{R}^{d_{\text{mono}}}$ (2)
where $W_{1}\in\mathbb{R}^{d_{\text{multi}}\times d_{\text{ffn}}}$ and $W_{2}\in\mathbb{R}^{d_{\text{ffn}}\times d_{\text{mono}}}$ are the weight matrices, $q$ is non-linearity function, and $d_{\text{mono}}$ represents the dimension of the target language representaiton. We set $d_{\text{ffn}}$ as $d_{\text{multi}}/4$, and the non-linearity function $q$ as SwiGLU (Shazeer, 2020). The output linear layer $g_{\text{mono}}$ then generates a word.
$\begin{split}f_{\text{mono}}(h)=g_{\text{mono}}(\text{FFN}(h))\in\mathbb{R}^{|\mathcal{V}_{\text{mono}}|}\end{split}$ (3)
@@ 3.2 Fine-tuning
$\underset{{f_{\text{mono}}}}{\max}~{}\mathcal{L}_{\text{MLE}}(p_{\text{mumo}},\mathbf{x})=\sum_{t=1}^{T}\log p_{\text{mumo}}(x_{t}|\mathbf{x}_{<t})$
@@ 3.3.1 Step 1: Top-$k$ Selection
Initially, we select top $k$ candidates based on the probability $p_{\text{mumo}}(x_{t}|\mathbf{x}_{<t})$. Given the fact that we do not modify the input embedding of the pre-trained model, we are unable to feed the predicted word if a word does not belong in $\mathcal{V}_{\text{multi}}$ during the subsequent iteration. Instead, we input the predicted word as the tokenization units of the pre-trained vocabulary. For example, let’s consider the Korean word “수소”, which corresponds to a sequence of two tokens (“수”, “소”) in $\mathcal{V}_{\text{multi}}$. If the Korean word “수소” is selected among the Top-$k$ candidates, we employ these two multilingual tokens.
@@ 3.3.2 Step 2: Verification
$\begin{split}\sigma(\mathbf{c}^{i})&=\frac{1}{l^{i}}\sum_{k=1}^{l^{i}}{\log p_{\text{multi}}(c_{t+k}^{i}|c_{<t+k}^{i},\mathbf{x}_{<t})}\end{split}$ (6)
@@ 4.1 Setup
We focus on two non-roman alphabetical languages: Korean and Japanese. These languages constitute over 0.05% of the pre-training corpus used in Llama-2 (Touvron et al., 2023). They are of particular interest due to their tokenization disparity when compared to English scripts (Ahia et al., 2023; Petrov et al., 2023). Notably, previous work (Bandarkar et al., 2023) has shown impressive results on natural language understanding tasks for these languages using Llama-2.
@@
Model
We utilize the Llama-2 13B model (Touvron et al., 2023) for all experiments. We observed some language alignment discrepancies between instructions and responses when using the Llama-2 13B chat model.111meta-llama/Llama-2-13b-chat To address the issue, we conduct multilingual instruction tuning (Muennighoff et al., 2022) for English, Korean, and Japanese languages using the ShareGPT and Alpaca (Chen et al., 2023c). This process improve the model’s fluency in each language (Muennighoff et al., 2022; Chen et al., 2023b).
Implementation of MuMo
To construct targeted monolingual vocabularies in MuMo Framework, we levergage the tokenizers from the off-the-shelf model, as shown in Table 3. We selected monolingual tokens by filtering vocabulary items based on the Unicode range of each monolingual script. Additionally, we excluded items from the selection if they were already present in the pre-trained vocabulary.
@@
We only train the Target Monolingual LM head $g_{\text{mono}}$ with the translated ShareGPT and Alpaca datasets (Chen et al., 2023c) in Korean, and Japanese. The training is done with 1500 steps, using 8 A100-40GB GPUs, with one batch consisting of 128 examples. We use the AdamW (Loshchilov and Hutter, 2019) optimizer with a learning rate of 0.001, weight decay of 0.01, and 150 steps of warm-up.
@@
Table 4: Comparative study of Language Model (LM) inference speed. The column labeled “Speed Up” represents the relative performance improvement in inference speed compared to the vanilla decoding method. The highest performance in each category is highlighted in Boldface, and the second highest score is underlined. All models use sampling-based decoding. MuMo outperforms the compared baselines in the inference speed. Detailed information about the generation hyperparameters, including those used for sampling-based decoding, can be found in Appendix A.

### v2

# Accelerating Multilingual Language Model for Excessively Tokenized Languages

@@ Abstract
Recent advancements in large language models (LLMs) have remarkably enhanced performances on a variety of tasks in multiple languages. However, tokenizers in LLMs trained primarily on English-centric corpora often overly fragment a text into character or Unicode-level tokens in non-Roman alphabetic languages, leading to inefficient text generation. We introduce a simple yet effective framework to accelerate text generation in such languages. Our approach involves employing a new language model head with a vocabulary set tailored to a specific target language for a pre-trained LLM. This is followed by fine-tuning the new head while incorporating a verification step to ensure the model’s performance is preserved. We show that this targeted fine-tuning, while freezing other model parameters, effectively reduces token fragmentation for the target language. Our extensive experiments demonstrate that the proposed framework increases the generation speed by a factor of 1.7 while maintaining the performance of pre-trained multilingual models on target monolingual tasks.
@@ 1 Introduction
Modern large language models (LLMs) (OpenAI, 2023; Touvron et al., 2023a; Antropic, 2023) have exhibited remarkable capabilities for a variety of tasks in multiple languages (Eloundou et al., 2023; Solaiman et al., 2023). Although these models are predominantly trained on English-centric data, they have shown a significant degree of multilingual proficiency (Bandarkar et al., 2023).
@@
Figure 1: Analysis of tokenization lengths and language distribution in pretraining corpus with percentage >=0.04% English script comprises 89.7% of the corpus and has an average token length of 29.6 in FLoRes-200. The languages using the Chinese, Japanese, and Korean (CJK) scripts have longer tokenization lengths compared to those using Latin and Cyrillic scripts. Our primary focus is on languages that are excessively tokenized by English-centric tokenizers.
@@
To address these challenges, previous studies Wang et al. (2019); Rust et al. (2021); Cui et al. (2023) have proposed replacing or augmenting the existing vocabulary of pre-trained multilingual models with language-specific vocabularies to more effectively encode monolingual text corpora. Specifically, Rust et al. (2021) improved mBERT Devlin et al. (2019) by replacing its tokenizer with a monolingual one and incorporating an additional 100,000 pre-training steps. On the other hand, Cui et al. (2023) enhanced Llama (Touvron et al., 2023a) by expanding the Chinese vocabulary and further pre-training it on a 120GB text corpus that includes Chinese texts. However, this approach requires an extensive pre-training phase with a substantial amount of data.
@@
In response to these challenges, our research introduces MuMo, accelerating Multilingual language models for a targeted Monolingual text generation, particularly in non-alphabetic languages. MuMo incorporates a new vocabulary of a target language into the output layer, also known as the Language Model (LM) head, and predicts the next token from this expanded vocabulary. This approach requires training only the extended portion of the output layer and specific layers of the feed-forward network. Importantly, MuMo eliminates the need for extensive text corpora or a draft model, requiring only a modest corpus of the target language, approximately 44M tokens in size. Empirical results across summarization, and translation tasks in Korean and Japanese demonstrate that MuMo significantly accelerates text generation, achieving over a 1.7-fold increase in speed without significantly compromising output quality.
2 Related Work
  Lang Word Multilingual Tokens Ko 서울 (“서”, “\0xec”, “\0xb8”, “\0x9a”) Ja {CJK*}UTF8min発売 ({CJK*}UTF8min“発”, “\0xe5”, “\0xa3”, “\0xb2”)
@@ 2 Related Work
Table 1: Examples of the tokenization results. These examples are preprocessed by the Llama tokenizer (Touvron et al., 2023b). The target monolingual word are excessively segmented into byte units, when a suitable match is not found in the multilingual vocabulary.
Tokenization Disparity
Subword tokenization, a common approach in LMs, is typically data-driven. Most of pre-trained tokenizers, which are often trained on predominantly English corpora, frequently result in excessive fragmentation of non-English scripts (Rust et al., 2021; Zhang et al., 2022). Ahia et al. (2023); Petrov et al. (2023) have found significant tokenization disparities across languages in popular LLMs (Xue et al., 2021, 2022; Scao et al., 2022; OpenAI, 2023). Our work endeavors to address the slowdown in inference that arises due to tokenization disparity in non-alphabetic languages.
Modifying Pre-trained Vocabulary
Previous works have explored the adaptation of pre-trained vocabularies or the addition of new tokens (Artetxe et al., 2020; Rust et al., 2021; Hong et al., 2021; Liu et al., 2023), these methods often necessitate extensive pre-training to integrate the new tokens effectively (Wang et al., 2019; Chau et al., 2020; Cui et al., 2023; Liu et al., 2023). In contrast, our MuMo framework sidesteps the need for fine-tuning the parameters of pre-trained models to preserve the original capabilities of the pre-trained language model. Efforts to select items of pre-trained embedding matrix have been made (Abdaoui et al., 2020; Domhan et al., 2022; Ushio et al., 2023), but these have not yielded significant speed up where the size of the embedding layer is relatively small  (Bogoychev et al., 2023).
Accelerating LLM Inference
The quest to accelerate inference in auto-regressive large language models (LLMs) has led to a variety of approaches. There has been a proliferation of systems specifically engineered for LLM inference (Yu et al., 2022; Sheng et al., 2023; Xiao et al., 2023). Our proposed methodology can be harmonically integrated with the aforementioned techniques. Speculative decoding (Leviathan et al., 2023; Chen et al., 2023a) have also been explored to increase inference velocity. However, the approach often relies on the assumption that a small model can maintain high fidelity when generating a series of multiple tokens. Moreover, acquiring a small yet competitive model may be tricky, especially in a multilingual setup (Conneau et al., 2020; Bandarkar et al., 2023). Our work distinguishes itself by specifically solving the inference inefficiency that arises from excessive fragmentation in the non-alphabetic context.
Parameter Efficient Cross-lingual Transfer Learning
The curse of multilinguality, which refers a trade-off between the language coverage and model capacity (Conneau et al., 2020), is a significant issue even in massively multilingual models, such as mBERT, XLM-R, and mT5 (Devlin et al., 2019; Conneau et al., 2020; Xue et al., 2021; Ansell et al., 2021). The problem has been mitigated through modular parameter-efficient adaptations of the multilingual models through lightweight adapters (Houlsby et al., 2019): additional trainable parameters inserted into the transformer layers of model (Pfeiffer et al., 2020; Üstün et al., 2020; Vidoni et al., 2020; Parović et al., 2022) for a target language. These techniques bear a resemblance to ours, in that they involve training partial parameters of a language model with a small amount of target language corpus. However, our goal is fundamentally different: we aim to accelerate the inference, whereas previous studies focus on improving the representational capability in target languages for multilingual models.
@@ 3 Proposed Framework
Figure 3: Illustration of a single-step prediction with MuMo. Initially, the MuMo LM Head $f_{\text{mumo}}$ selects the top 6 candidates. Then, the pre-trained multilingual model verifies the feasibility of the candidates. Among the modules in MuMo, the Target Monolingual LM head (the Korean LM Head in the figure) is only trained.
We propose a framework named MuMo to accelerate the inference speed of a pre-trained multilingual LM for a non-alphabetic monolingual language via a given small monolingual dataset. In the section, we introduce 1) the model architecture, 2) the fine-tuning process on a small targeted language dataset, and 3) the inference process of the proposed framework.
@@ 3.1 Model Architecture
We consider a setting in which a pre-trained multilingual model $f_{\text{multi}}$ is given. The model consists of 1) Transformer layers that consist of attention and feed-forward network, and 2) an output embedding layer called language model (LM) head. We denote $\mathcal{V}_{\text{multi}}$ as the multilingual vocabulary set of the model objective, as $\mathcal{L}_{\text{MLE}}(p_{\text{multi}},\mathbf{x})=\sum_{t=1}^{|\mathbf{x}|}\log p_{\text{multi}}(x_{t}|\mathbf{x}_{<t})$,
@@
The primary concept involves modifying pre-trained representations to predict a singular token unit within a target monolingual vocabulary $\mathcal{V}_{\text{mono}}$. The Target Monolingual LM head $f_{\text{mono}}$ projects the hidden representation $h$, which is composed of two main components: a feed-forward network (FFN) and an output linear layer, represented as $g_{\text{mono}}:\mathbb{R}^{d_{{}_{\text{mono}}}}\rightarrow\mathbb{R}^{|\mathcal{V}_{\text{mono}}|}$:
$\text{FFN}(h)=q(W_{1}^{\top}h)W_{2}\in\mathbb{R}^{d_{\text{mono}}},$ (1)
where $W_{1}\in\mathbb{R}^{d_{\text{multi}}\times d_{\text{ffn}}}$ and $W_{2}\in\mathbb{R}^{d_{\text{ffn}}\times d_{\text{mono}}}$ are the weight matrices, $q$ is non-linearity function, and $d_{\text{mono}}$ represents the dimension of the target language representaiton. We set $d_{\text{ffn}}$ as $d_{\text{multi}}/4$, and the non-linearity function $q$ as SwiGLU (Shazeer, 2020). The output linear layer $g_{\text{mono}}$ then generates a subword token:
$\begin{split}f_{\text{mono}}(h)=g_{\text{mono}}(\text{FFN}(h))\in\mathbb{R}^{|\mathcal{V}_{\text{mono}}|}.\end{split}$ (2)
@@ 3.2 Fine-tuning
[…]
@@ 3.3.1 Step 1: Top-$k$ Selection
Initially, we select top-$k$ candidates based on the probability $p_{\text{mumo}}(x_{t}|\mathbf{x}_{<t})$. We set $k$ as 10 for all experiments. Given the fact that we do not modify the input embedding of the pre-trained model, we are unable to feed the predicted word if a word does not belong in $\mathcal{V}_{\text{multi}}$ during the subsequent iteration. Instead, we input the predicted word as the tokenization units of the pre-trained vocabulary. For example, let’s consider the Korean word “수소”, which corresponds to a sequence of two tokens (“수”, “소”) in $\mathcal{V}_{\text{multi}}$. If the Korean word “수소” is selected among the Top-$k$ candidates, we employ these two multilingual tokens.
@@ 3.3.2 Step 2: Verification
$\begin{split}\sigma(\mathbf{c}^{i})&=\frac{1}{l^{i}}\sum_{k=1}^{l^{i}}{\log p_{\text{multi}}(c_{t+k}^{i}|c_{<t+k}^{i},\mathbf{x}_{<t})},\end{split}$ (5)
@@ 4.1 Setup
As a case study, we focus on two non-roman alphabetical languages: Korean and Japanese. Since we aimed to utilize a pre-trained model with a reasonable level of effectiveness in the target language, it is essential that the language is explicitly mentioned as being trained within the pre-training corpus. In this context, we considered languages included in the Llama-2 (Touvron et al., 2023b) pre-training corpus. Moreover, the chosen language needed to exhibit the excessive fragmentation problem (Ahia et al., 2023; Petrov et al., 2023) by the English-centric pre-trained tokenizer. (See the Figure 1) This criterion led to the exclusion of most European languages such as French, German, and Portuguese. Finally, we conduct a study on multiple tasks, necessitating the existence of an instruction dataset for the target language. Due to these considerations, we only implement the experiment in Korean and Japanese.
Model
We utilize the Llama-2 13B model (Touvron et al., 2023b) for all experiments. We observed some language alignment discrepancies between instructions and responses when using the Llama-2 13B chat model.111meta-llama/Llama-2-13b-chat To address the issue, we conduct multilingual instruction tuning (Muennighoff et al., 2022) for English, Korean, and Japanese languages using the ShareGPT and Alpaca (Chen et al., 2023c). This process improve the model’s fluency in each language (Muennighoff et al., 2022; Chen et al., 2023b). We also report our results test on Llama-1 13B (Touvron et al., 2023a) in Appendix.
Implementation of MuMo
To construct targeted monolingual vocabularies in MuMo Framework, we levergage the tokenizers from the off-the-shelf model, as shown in Table 2. We selected monolingual tokens by filtering vocabulary items based on the Unicode range of each monolingual script. Additionally, we excluded items from the selection if they were already present in the pre-trained vocabulary. In terms of the preprocessing algorithm, we employ a forward maximum matching strategy to identify words in a target language vocabulary. This strategy identifies the longest sequence of tokens that aligns with a word in the target language vocabulary.
@@
We only train the Target Monolingual LM head $g_{\text{mono}}$ with the translated ShareGPT and Alpaca datasets (Chen et al., 2023c) in Korean, and Japanese. The training is done with 1500 steps with one batch consisting of 128 examples. We use the AdamW (Loshchilov and Hutter, 2019) optimizer with a learning rate of 0.001, weight decay of 0.01, and 150 steps of warm-up.
@@
Table 3: Comparative study of Language Model (LM) inference speed. The column labeled “Speed Up” represents the relative performance improvement in inference speed compared to the vanilla decoding method. The highest performance in each category is highlighted in Boldface, and the second highest score is underlined. All models use sampling-based decoding. MuMo outperforms the compared baselines in the inference speed. Detailed information about the generation hyperparameters, including those used for sampling-based decoding, can be found in Appendix D.